OpenStudy (anonymous):

solve for x in x-(x+11)^1/2=1 the ^1/2 means square root in the question solve for x in x-(x+11)^1/2=1 the ^1/2 means square root in the question @Mathematics

OpenStudy (agreene):

x=5

OpenStudy (anonymous):

how is it 5?

OpenStudy (agreene):

x-(x+11)^(1/2)=1 -(x+11)^(1/2)=1-x (x+11)^(1/2)=x-1 x+11=(x-1)^2 x+11=x^2-2x+1 -x^2+3x+10=0 x^2-3x-10=0 x^2-3x=10 x^2-3x+9/4=49/4 (x-3/2)^2=49/4 |x-3/2|=7/2 x-3/2 = +- 7/2 x=-2 or x=5 plug in and see -2 doesnt work, 5 does thus x=5 QED